# A contramodule generalization of Neeman’s flat and projective module theorem

### v1

@@ 7. Projective and Flat Constructions of Contraderived Category
    $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ belongs to the closure of the class of all contractible complexes of flat $\mathfrak{R}$-contramodules under extensions and directed colimits;
@@
    $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is an $\aleph_{1}$-directed colimit of totalizations of short exact sequences of complexes of countably presentable flat $\mathfrak{R}$-contramodules.
@@
The equivalences (iic) $\Longleftrightarrow$ (iiic) $\Longleftrightarrow$ (ivc) are provided by [26, Theorem 13.2]. ∎
Finally, we deduce a corollary about an equivalence of the “flat” and “projective” constructions of the contraderived category $\mathsf{D}^{\mathsf{bctr}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}})$. It should be compared with [32, Theorems 4.12 and 5.7]. As usual, we denote by $\mathsf{D}(\mathsf{E})$ the (conventional) unbounded derived category of an exact category $\mathsf{E}$.
@@
$\mathsf{Hot}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{proj}})\simeq\mathsf{D}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})\simeq\mathsf{D}^{\mathsf{bctr}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}).$
@@
The assertion that that the functor $\mathsf{D}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})\longrightarrow\mathsf{D}^{\mathsf{bctr}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}})$ is a triangulated equivalence is a corollary of Theorem 7.1 (ic) $\Leftrightarrow$ (iic). Consider a complex of flat left $\mathfrak{R}$-contramodules $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$. Then, as a particular case of the previous paragraph, there exists a complex of projective left $\mathfrak{R}$-contramodules $\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ together with a morphism of complexes $\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ with a contraacyclic cone. The cone, being a contraacyclic complex of flat $\mathfrak{R}$-contramodules, is consequently an acyclic complex in the exact category $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$. So $\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is an isomorphism in $\mathsf{D}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})$. Furthermore, every complex of projective $\mathfrak{R}$-contramodules that is acyclic in $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$ is contractible. The desired conclusion easily follows. ∎
Question 7.3.
It would be very interesting to extend the result of Theorem 7.1 to complete, separated topological rings $\mathfrak{R}$ with a countable base of neighborhoods of zero consisting of open right ideals. Is such a generalization true? See [26, Questions 9.8, 11.2, and 13.3] for further discussion.

### v3

@@ 11. Five Constructions of the Contraderived Category
    $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ can be obtained from contractible complexes of flat $\mathfrak{R}$-contramodules using extensions and directed colimits;
@@
    $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is an $\aleph_{1}$-directed colimit of total complexes of short exact sequences of complexes of countably presentable flat $\mathfrak{R}$-contramodules;
  5. (vc)
    for every complex of cotorsion left $\mathfrak{R}$-contramodules $\mathfrak{C}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$, any morphism of complexes of $\mathfrak{R}$-contramodules $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow\mathfrak{C}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is homotopic to zero;
  6. (vic)
    for every complex of flat cotorsion left $\mathfrak{R}$-contramodules $\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$, any morphism of complexes of $\mathfrak{R}$-contramodules $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is homotopic to zero.
@@
The equivalences (iic) $\Longleftrightarrow$ (iiic) $\Longleftrightarrow$ (ivc) are provided by [34, Theorem 13.2].
The equivalence (iic) $\Longleftrightarrow$ (vc) is Corollary 8.4.
The equivalence (iic) $\Longleftrightarrow$ (vic) is Corollary 9.1. ∎
Finally, we deduce a corollary about an equivalence of the “flat”, “projective”, “flat cotorsion”, “Verdier quotient by contraacyclic”, and “Verdier quotient of cotorsion by contraacyclic” constructions of the contraderived category $\mathsf{D}^{\mathsf{bctr}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}})$. It should be compared with [43, Theorems 4.12, 4.13, and 5.7].
@@
$\mathsf{Hot}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{proj}})\simeq\mathsf{D}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})=\mathsf{D}^{\mathsf{bctr}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})\simeq\mathsf{D}^{\mathsf{bctr}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}).$
Furthermore, the inclusions of additive/exact/abelian categories $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}^{\mathsf{cot}}\longrightarrow\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$ and $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}^{\mathsf{cot}}\longrightarrow\mathfrak{R}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}}\longrightarrow\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$ induce equivalences of triangulated categories
$\mathsf{Hot}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}^{\mathsf{cot}})\simeq\mathsf{D}^{\mathsf{bco}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})=\mathsf{D}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})$
and
$\mathsf{Hot}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}^{\mathsf{cot}})\simeq\mathsf{D}^{\mathsf{bctr}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}})\simeq\mathsf{D}^{\mathsf{bctr}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}).$
@@
The assertion that the functor $\mathsf{D}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{proj}})\longrightarrow\mathsf{D}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})$ is a triangulated equivalence is a corollary of Theorem 11.1 (ic) $\Leftrightarrow$ (iic). We apply Lemma 2.9. Consider a complex of flat left $\mathfrak{R}$-contramodules $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$. Then, as a particular case of the previous paragraph, there exists a complex of projective left $\mathfrak{R}$-contramodules $\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ together with a morphism of complexes $\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ with a contraacyclic cone. The cone, being a contraacyclic complex of flat $\mathfrak{R}$-contramodules, is consequently an acyclic complex in the exact category $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$. So $\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is an isomorphism in $\mathsf{D}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})$. Furthermore, every complex of projective $\mathfrak{R}$-contramodules that is acyclic in $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$ is contractible, again by Theorem 11.1 (ic) $\Leftrightarrow$ (iic) or by [34, Proposition 12.1]. The desired conclusion follows.
One has $\mathsf{D}^{\mathsf{bctr}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})=\mathsf{D}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})=\mathsf{D}^{\mathsf{bco}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})$ by Theorem 11.1 (ic) $\Leftrightarrow$ (iic) $\Leftrightarrow$ (vic). The triangulated equivalence $\mathsf{Hot}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}^{\mathsf{cot}})\simeq\mathsf{D}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})$ is the result of Corollary 9.2. The triangulated equivalence $\mathsf{Hot}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}^{\mathsf{cot}})\simeq\mathsf{D}^{\mathsf{bctr}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}})$ is the result of Corollary 10.3.
Finally, the triangulated functor $\mathsf{D}^{\mathsf{bctr}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}})\longrightarrow\mathsf{D}^{\mathsf{bctr}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}})$ induced by the inclusion of exact/abelian categories $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}}\longrightarrow\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$ is well-defined by Corollary 10.2. This functor is a triangulated equivalence, since both the composition $\mathsf{Hot}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}^{\mathsf{cot}})\longrightarrow\mathsf{D}^{\mathsf{bctr}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})\longrightarrow\mathsf{D}^{\mathsf{bctr}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}})$ and the functor $\mathsf{Hot}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}^{\mathsf{cot}})\longrightarrow\mathsf{D}^{\mathsf{bctr}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}})$ are triangulated equivalences. ∎
Question 11.3.
It would be very interesting to extend the result of Theorem 11.1 (or some parts of it) to complete, separated topological rings $\mathfrak{R}$ with a countable base of neighborhoods of zero consisting of open right ideals. Is such a generalization true? See [34, Questions 9.8, 11.2, and 13.3] for further discussion.